Suppose that the life span of a certain automobile tire is normally distributed with mu equals 23,000 miles and sigmaequals2500 miles.

​(a) Find the probability that a tire will last between 28,000 and 30,500 miles. ​

(b) Find the probability that a tire will last more than 29,000 miles.
Answer #1

Please don't hesitate to give a "thumbs up" for the answer in case you're satisfied with the answer

We have the parameter of normal distribution :

Mu = 23000

Stdev = 2500

a. P(28000<X<30500) = ?

= P(28000-23000/2500 <Z< 30500-23000/2500)

= P(2<Z<3)

= .0214

b.

P(X>29000)

= P(Z> 29000-23000 / 2500)

= P(Z> 6000/2500)

= 0.0082